A body A of mass m1 = 5 kg moving along a directed axis in the positive sense with a velocity of 15 m/s collides head-on with a body B of mass m2 = 8 kg moving in the negative sense with a velocity of 4 m/s. If the velocity of m1 after the collision is 4m/s in the negative sense. Apply your scientific knowledge and understanding of impulse and momentum theorem to find the final velocity of m2
1
Expert's answer
2020-11-17T11:26:12-0500

Solution

We are applying momentim conservation

"m_1u_1+m_2u_2=m_1v_1+m_2v_2"

Where

u1and u2 are initial velocity

And v1 and v2 are final velocity.

Using all data in question

"5\times15-8\times4=-5\times4+8\times v_2\\v_2=7.8m/s"